[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Sep 25, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Sep 25 09:42:12 UTC 2017


HSU Shift Log for Mon Sep 25 2017 05:42:05 GMT-0400 (EDT)

Shift: A

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: EPAC ATCF hung up for a while near 0900 UTC. I'm currently  
re-starting the unit.

N-AWIPS issues: EPAC N-AWIPS terminal SAWFISH also hung up near 0900 UTC.  
I'm currently trying to re-start that unit.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170925/bb2dacf3/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list